Consider a bond with a par value of £1,000, a coupon rate of 8%, a maturity of 20 years and one coupon instalment per year. The market interest rate is currently 10%. Suppose you buy the bond today with the view to sell it in a year’s time. What would be your return on the bond if interest rate were to rise to 15% in a year’s time? 


1
Expert's answer
2021-02-09T07:03:39-0500

Solutions:

Derive the annual interest rate:

Annual interest rate ="Par \;value\times coupon \;rate"

                                = "\pounds1000\times 8\% = \pounds80"


Yield To Maturity (YTM) interest rate = 15%


Return on the bond = "\frac{C}{N}[\frac{1 - ((1+\frac{R}{N}))^{2} (-N\times T))) ]}{\frac{R}{N}}]+[\frac{F}{((1+\frac{R}{N}))^{2}(-N\times T)))} ]"

Where:

C = Annual interest = £80

N = Number of payments per year = 1

R = YTM = 0.15

F = Par value = £1000

T = Number of years until maturity = 20 years


Plug the figures into the formula:

"\frac{80}{1}[\frac{1 - ((1+\frac{0.15}{1}))^{2} (-1\times 20))) ]}{\frac{0.15}{1}}]+[\frac{1000}{((1+\frac{0.15}{1}))^{2}(-1\times 20)))} ]"


"\frac{80}{1}[1-(\frac{(1.15^{2} -20)}{0.15}]+[\frac{1000}{1.15^{2} 20)}]"


"(80) (\frac{1-0.0611}{0.15} )+(\frac{1000}{16.3665} )"

"(80) (\frac{0.9389}{0.15} )+(61 )"


"500.75+61 = \pounds561.75"


Return on the bond = £561.75